OpenStudy (anonymous):

The preimeter of a traffic sign must be 136 inches. Also its length is 14 inches longer than its width. Find the dimensions.

OpenStudy (anonymous):

2L+2w = 136 l = w +14 2(w+14) + 2w = 136 2w + 28 + 2w = 136 4w = 136 - 28 4w = 108 w =27 w + 14 = 41 l = 41